Jefferson Lab Weekly Briefs May 18, 2011
Jefferson Lab Weekly Briefs
May 18, 2011
Last week, Hall A completed the short-range correlations experiment on Friday. Hall staffers are entering the six-month shutdown with a big list of work items to prepare for the g2p/GEp experiments.
The Q-weak experiment in Hall C completed a six-month commissioning and initial data taking run, acquiring an estimated 30 percent of the proposed data. The reliable target performance and excellent accelerator operations allowed the experiment to commission the detectors and Compton polarimeter, begin data taking, and learn lessons for the next run. Q-weak will run for another six months, starting in November.
The CEBAF accelerator operations crew terminated beam operations for FY11 on May 13. The CEBAF acclerator has been operating continuously since the beginning of the fiscal year in October, with only the short winter break at the end of December. The final week of operations was occupied with beam delivery to Halls A and C and many beam studies. The beam studies targeted the even arc dipole field gradients; these dipoles will be modified during the six-month shutdown, and these measurements will help resolve any difficulties with beam restoration after the shutdown. Beam studies were also used to characterize the 2L21 cryomodule for comparison to the new cryomodules being characterized in the Test Lab. At 7 a.m. on May 13, beam was terminated and the shutdown began. Thanks to everyone who participated in FY11 CEBAF operations!
The FEL group continued multi-kilowatt running into last weekend and completed the high-power run on the oxide-coated outcoupler. This week, the team successfully completed the planned set of tasks for the Air Force, while running pulsed laser deposition experiments on swing and partial owl shifts. The FEL has now shut down for installation of a new low-conductivity water system.
Theoretical and Computational Physics
A new theory paper <arXiv:1105.2764> explores the possibility of incorporating the Roper resonance into chiral perturbation theory (ChPT), a low-energy effective theory for QCD that in its quintessential form includes only the pion and the nucleon. Although it is expected that the Roper would improve the long range interactions in the π N system, the fact that the Roper-nucleon mass splitting is large may jeopardize the convergence of the low momentum expansion in ChPT. This article investigates the effects of the Roper on low energy π N scattering and discusses the profitability of including the Roper as an explicit degree of freedom in ChPT.
The alignment group (Metrology) is surveying the west and east arc dipoles and many of the quadrupole magnets in support of the six-month shutdown. The magnets' current positions must be found (to better than 0.25 millimeter). Over the years (originally installed 1992-1994), the tunnel has shifted slightly, and the magnets have also shifted due to this movement and probably also due to cycles in temperature of the supports. The magnets will be refurbished and have ‘H’ steel added. Following this procedure, operations personnel want the "as-found" magnet positions re-established to reduce the amount of changes they will have to make to thread the beam through the magnets on its path to the end stations.
Additionally, alignment is carrying out a precise survey in the northeast corner of the accelerator, where there is excavation to extend the north linac to the new Hall D tagger building. This activity will be used to monitor any movements that may occur in the tunnel and the magnets due to the removal of the earth covering the existing tunnel and the actual construction. If any movement is detected after construction, the magnets will have to be re-aligned in this area.
Precision levels are also being done throughout the accelerator to update the existing alignment control network, and two cryomodules will be aligned for test purposes in the next few days.
The majority of the 'H' steel has arrived onsite. The ‘H’ steel will be added to the dipoles pulled from the tunnel. The magnet group has completed refurbishment trials for the BB (two-meter) and BA (three-meter) dipole magnets. The baseline magnetic measurements have been completed for the BB and BA magnets.
The vacuum chambers and synchrotron radiation coils were installed on 31 four-meter XP dipole magnets. Stretch wire measurements were completed on these dipoles. Twenty eight of the XPs will be installed in the tunnel during the six-month down.
Other tasks completed by the magnet group were the completion of the six month refurbishment schedule, building new fixtures for the stretch wire measurement tables and upgrading the hall probe holding fixture for detail grid measurements.
Environment, Safety, Health and Quality
Critters, Bugs and Reptiles On the Prowl at JLab
If you should come across any wildlife in circumstances that cause you concern, alert Mike Lewellen, x7169, as soon as feasible and follow up by putting in an FM&L Work Request. Do not confront or corner wildlife – FM&L retains the services of a pest control company and can safely deal with your concern.
If you are stung or bitten, report to Occupational Medicine. Prompt medical care can help prevent infection and complications that could result in unnecessary pain and lost work time.
Computing and Networking Infrastructure
Maintenance Day Scheduled for May 24
Mont on Big Numbers
Library Is A'Changing
Book Swap Lunch Party Planned for Friday, May 27
Tracking Thomas at JLab
Congratulations to Kevin Jordan, David Fazenbaker, Jason Willoughby and Debra Brand who were the first people to correctly identify the May 11 location. Honorable mentions go out to Elena Long, Keith Royston, Gina Dixon, Sue Witherspoon, Harry Fanning, Mike Koch, Charlie Garrison, Stephanie Vermeire, Dick Owen, Stephen Smith, Mike Lewellen and Bill Merz. Check out the Tracking Thomas webpage for a better view of his last location and this week's new mystery photo.
JLab Calendar of Events
May 17-20: NSTAR 2011